### v1

@@ II.3 Decuplet saturation approximation
Here $\Delta$ denotes the average decuplet–octet baryon mass splitting, and we take $\Delta=203$ MeV, which is evaluated using the masses of the baryons considered, namely, the proton, neutron, $\Delta(1232)$, $\Xi^{0}$, $\Xi^{-}$, and $\Xi(1530)$ [20]. We employ the large-$N_{c}$ value for the LEC $C$, namely, $C=3g_{A}/4$} [12, 21, 22]. Then, under the DSA, the LECs $G_{e}^{(\sigma)}$ and $G_{f}^{(\sigma)}$, which enter $U_{\mathrm{OPE}}^{(\sigma)}$, become independent of $\sigma$;
@@ II.4 Correlation function
where $F_{L}$ is the regular Coulomb wave function with the orbital angular momentum $L$ and $q$ is the relative momentum between $d$ and $\Xi^{-}$. The source function $\mathcal{S}$ used here is the same as that employed in Ref. [11];
@@
$V_{\mathrm{C}}(R)=\begin{dcases}-\frac{e^{2}}{2R_{\mathrm{C}}}\left(3-\frac{R^{2}}{R_{\mathrm{C}}^{2}}\right)&(R\leq R_{\mathrm{C}}),\\ -\frac{e^{2}}{R_{\mathrm{C}}}&(R>R_{\mathrm{C}}),\end{dcases}$ (77)
@@ III.1 Comparison of potentials
Figure 6(a) shows $U_{\mathrm{TPE}}^{(u_{3})}$, $U_{\mathrm{OPE}}^{(\sigma)}$, and $U_{\mathrm{ct}}^{(\sigma)}$ as a function of $R$, and is depicted by the solid, dashed, and dotted lines, respectively. Note that $U_{\mathrm{TPE}}^{(u_{1})}$ defined by Eq. (47) becomes zero under the DSA. As a representative case, we fix both $G^{\prime}=1$} and $H^{\prime}=1$}, with the cutoff $R_{0}=1.0$} fm. One sees that the contact term can have the largest magnitude at $R\lesssim 1.0$} fm. Beyond this range, the TPE and OPE terms become dominant although their absolute magnitude is much smaller compared to that of the contact term near the origin.
In Fig. (6)(b), we compare $U_{\mathrm{3BF}}^{(\sigma)}$ with $U_{\mathrm{2BF}}^{(\sigma)}$. In this case, we select two sets of the LECs, one giving the most attractive and the other the most repulsive potential, in order to maximize the variation in $U_{\mathrm{3BF}}^{(\sigma)}$. These parameter sets correspond to the solid and dashed lines, with $(G^{\prime},H^{\prime})=(1.0,-1.0)$} and $(G^{\prime},H^{\prime})=(-1.0,1.0)$}, respectively. The dotted and dash-dotted lines are $U_{\mathrm{2BF}}^{(\sigma)}$ with $\sigma=1/2$ and $3/2$, respectively. In both sets of the LECs, $U_{\mathrm{3BF}}^{(\sigma)}$ are found to be weaker in magnitude and shorter in range compared to $U_{\mathrm{2BF}}^{(\sigma)}$.
@@ III.2 Effect of $\Xi NN$ three-baryon force on correlation function
[…]
@@ III.3 Variation of parameters
To conclude this section, we consider whether the loosely-bound nature of the deuteron can explain the small effect of the $\Xi NN$ 3BF on $C_{d\Xi^{-}}$. To examine this, we artificially change the deuteron binding energy $E_{\mathrm{BE}}$ by modulating the strength of the Argonne $V4^{\prime}$ potential. Figures 10(a), (b), and (c) show $U_{\mathrm{2BF}}^{(\sigma)}$ and $U_{\mathrm{3BF}}^{(\sigma)}$ computed with $E_{\mathrm{BE}}/\mathrm{nucleon}=1.12$$2.54$, and $6.98$ MeV, corresponding to that of the realistic deuteron, triton and $\alpha$ particle, respectively. Even though $E_{\mathrm{BE}}$ is increased, the contribution of $U_{\mathrm{3BF}}^{(\sigma)}$ relative to $U_{\mathrm{2BF}}^{(\sigma)}$ does not change significantly. Figures 10(d), (e), and (f) display $C_{d\Xi^{-}}$ computed with the potentials in Figs. 10(a), (b), and (c), respectively. With increasing $E_{\mathrm{BE}}$, the magnitudes of the potentials grow, enhancing $C_{d\Xi^{-}}$. This enhancement mainly originates from $U_{\mathrm{2BF}}^{(\sigma)}$ rather than $U_{\mathrm{3BF}}^{(\sigma)}$. Indeed, even for the strongly-bound case in Fig. (10)(f), the effect of the $\Xi NN$ 3BF remains below $8\%$. Therefore, we conclude that the weak binding of the deuteron is not the primary reason for the small effect of the 3BF on the correlation function.
IV Summary and perspectives
@@
In conclusion, given the limited sensitivity of the correlation function to such short-range forces, it may be more promising to explore their impact on other observables, such as deuteron–hyperon scattering cross sections. Experimentally, nucleon–hyperon scattering became available recently [27, 28, 29]. Therefore, these complementary approaches could provide further insight into the role of 3BFs in many-hyperon–nucleon systems.
@@ Acknowledgements.
The authors thank T. Hyodo, A. Jinno, S. Ogawa, and Y. Chazono for valuable comments and discussions. This work was supported by JST ERATO Grant Number JPMJER2304, as well as JSPS KAKENHI Grant Numbers JP21K13919 and JP23KK0250.
@@ Appendix A SU(3) coefficients of vertices relevant to $\Xi NN$ three-baryon force
Table 1: SU(3) coefficients $N_{BB\phi}$ relevant to the $\Xi NN$-TPE and OPE potentials. The axial vector coupling coefficients $g_{A}$ and $g_{B}$ are defined in Sect. II.1.2.
@@
Table 4: Lists of the contact coefficients $N_{\begin{subarray}{c}BBB\\ BBB\end{subarray}}^{a}$ defined by $N_{\begin{subarray}{c}BBB\\ BBB\end{subarray}}^{a}=\sum_{f=1}^{11}\tilde{t}^{f,a}N_{\begin{subarray}{c}BBB\\ BBB\end{subarray}}^{f,a}$, which enter the $\Xi NN$-contact potentials as in Eqs. (33). The coefficients are all written in terms of the LECs $C_{i}$ introduced in Sect. II.1.4.
@@ B.1 Central components of $\Xi NN$ potential
$X_{123(0)}^{456}$ $=\dfrac{g_{B}d}{6f_{\pi}^{4}}\dfrac{q_{1}^{2}}{q_{1}^{2}+m_{\pi}^{2}}\left[(\bm{\tau}_{2}-\bm{\tau}_{3})\cdot\bm{\tau}_{1}\,(\bm{\sigma}_{2}-\bm{\sigma}_{3})\cdot\bm{\sigma}_{1}+(\bm{\tau}_{1}\times\bm{\tau}_{2})\cdot\bm{\tau}_{3}\,(\bm{\sigma}_{1}\times\bm{\sigma}_{2})\cdot\bm{\sigma}_{3}\right],$ (84)
$X_{231(0)}^{564}$ $=\dfrac{g_{A}}{6f_{\pi}^{4}}\dfrac{q_{2}^{2}}{q_{2}^{2}+m_{\pi}^{2}}\left[(e_{1}\bm{\sigma}_{3}+e_{2}\bm{\sigma}_{1})\cdot\bm{\sigma}_{2}\,\bm{\tau}_{2}\cdot\bm{\tau}_{3}+(e_{3}\bm{\sigma}_{3}+e_{4}\bm{\sigma}_{1})\cdot\bm{\sigma}_{2}\,\bm{\tau}_{1}\cdot\bm{\tau}_{2}-e_{5}(\bm{\tau}_{1}\times\bm{\tau}_{2})\cdot\bm{\tau}_{3}\,(\bm{\sigma}_{1}\times\bm{\sigma}_{2})\cdot\bm{\sigma}_{3}\right],$ (85)
$X_{312(0)}^{645}$ $=\dfrac{g_{A}}{6f_{\pi}^{4}}\dfrac{q_{3}^{2}}{q_{3}^{2}+m_{\pi}^{2}}\left[(e_{4}\bm{\sigma}_{1}+e_{3}\bm{\sigma}_{2})\cdot\bm{\sigma}_{3}\,\bm{\tau}_{3}\cdot\bm{\tau}_{1}+(e_{2}\bm{\sigma}_{1}+e_{1}\bm{\sigma}_{2})\cdot\bm{\sigma}_{3}\,\bm{\tau}_{2}\cdot\bm{\tau}_{3}-e_{5}(\bm{\tau}_{1}\times\bm{\tau}_{2})\cdot\bm{\tau}_{3}\,(\bm{\sigma}_{1}\times\bm{\sigma}_{2})\cdot\bm{\sigma}_{3}\right],$ (86)
$X_{312(0)}^{564}$ $=\dfrac{g_{A}}{6f_{\pi}^{4}}\dfrac{q_{3}^{2}}{q_{3}^{2}+m_{\pi}^{2}}\left[\{f_{1}\bm{\sigma}_{1}+f_{2}\bm{\sigma}_{2}+f_{3}i(\bm{\sigma}_{1}\times\bm{\sigma}_{2})\}\cdot\bm{\sigma}_{3}\,\bm{\tau}_{3}\cdot\bm{\tau}_{1}+\{f_{2}\bm{\sigma}_{1}+f_{1}\bm{\sigma}_{2}+f_{3}i(\bm{\sigma}_{1}\times\bm{\sigma}_{2})\}\cdot\bm{\sigma}_{3}\,\bm{\tau}_{2}\cdot\bm{\tau}_{3}\right$.
@@
$X_{231(0)}^{645}$ $=\dfrac{g_{A}}{6f_{\pi}^{4}}\dfrac{q_{2}^{2}}{q_{2}^{2}+m_{\pi}^{2}}\left[\{f_{1}\bm{\sigma}_{3}+f_{2}\bm{\sigma}_{1}-f_{3}i(\bm{\sigma}_{3}\times\bm{\sigma}_{1})\}\cdot\bm{\sigma}_{2}\,\bm{\tau}_{2}\cdot\bm{\tau}_{3}+\{f_{2}\bm{\sigma}_{3}+f_{1}\bm{\sigma}_{1}-f_{3}i(\bm{\sigma}_{3}\times\bm{\sigma}_{1})\}\cdot\bm{\sigma}_{2}\,\bm{\tau}_{1}\cdot\bm{\tau}_{2}\right$.
@@ B.2 Matrix elements of spin–isospin operators
with $\mu=\pm 1$. In Eq. (90), $\sigma$ and $m_{\sigma}$ represent the total spin of the system and its third component, respectively. The isospin and its third component of the system are fixed as $1/2$ and 1/2-1/2, respectively. Using these states, we evaluate the matrix elements of all spin and isospin operators involved in the potentials given in Sect. B.1.
Hereafter we redefine the particle indices $\alpha$ as $\Xi\to 1$, $p\to 2$, and $n\to 3$. The matrix elements of the spin operators, $\bm{\sigma}_{1}\cdot\bm{\sigma}_{2}$ and $\bm{\sigma}_{1}\cdot\bm{\sigma}_{3}$, depend on $\sigma$;
$\Braket{\Upsilon_{\sigma m_{\sigma}}|\bm{\sigma}_{1}\cdot\bm{\sigma}_{2}|\Upsilon_{\sigma m_{\sigma}}}$ $=\Braket{\Upsilon_{\sigma m_{\sigma}}|\bm{\sigma}_{3}\cdot\bm{\sigma}_{1}|\Upsilon_{\sigma m_{\sigma}}}=\begin{dcases}-2&(\sigma=1/2),\\ 1&(\sigma=3/2).\end{dcases}$ (93)
The other spin operators yield $\sigma$-independent expressions;
@@ B.3 Fourier transform and multipole expansion
where the functions $Y$ and $Z$ are defined by Eqs. (54) and (56), respectively. The arguments of the spherical harmonics $Y_{\lambda\mu}$ ($\lambda\leq\mu\leq\lambda$) are defined by $\hat{\bm{r}}_{i}=\bm{r}_{i}/r_{i}$. To carry out the above calculations, the following relation is used [31];
@@
$W^{(\sigma)}(\bm{r},\bm{R})$ $=W_{\mathrm{TPE}}(\bm{r},\bm{R})+W_{\mathrm{OPE}}^{(\sigma)}(\bm{r},\bm{R})+W_{\mathrm{ct}}^{(\sigma)}(\bm{r},\bm{R})$ (103)
$W_{\mathrm{TPE}}(\bm{r},\bm{R})$ $=-\frac{\sqrt{3}g_{A}^{2}m_{\pi}^{6}u_{1}}{4\pi f_{\pi}^{4}}Z_{1}(m_{\pi}r_{2})Z_{1}(m_{\pi}r_{3})\left[Y_{1}\left(\hat{\bm{r}}_{2}\right)\otimes Y_{1}\left(\hat{\bm{r}}_{3}\right)\right]_{00}$
$-\dfrac{g_{A}^{2}m_{\pi}^{6}u_{3}}{12\pi f_{\pi}^{2}}\left[\dfrac{1}{8\pi}Y(m_{\pi}r_{2})Y(m_{\pi}r_{3})+\dfrac{1}{\sqrt{5}}Z(m_{\pi}r_{2})Z(m_{\pi}r_{3})\left[Y_{2}\left(\hat{\bm{r}}_{2}\right)\otimes Y_{3}\left(\hat{\bm{r}}_{2}\right)\right]_{00}\right],$ (104)
@@
where $\bm{r}_{2}$ and $\bm{r}_{3}$ can be expressed as functions of $\bm{r}$ and $\bm{R}$ by Eqs. (43) and (44). The function $Z_{1}$, as well as the constants $G_{e}^{(\sigma)}$, $G_{f}^{(\sigma)}$, and $H^{(\sigma)}$, are all defined in Sect. II.2.
Then, by applying the regularization described in Sect. II.2 to $W^{(\sigma)}$, we can perform the integration over $\bm{r}$ and $\hat{\bm{R}}$ in Eq. (45), for which the multipole expansion is applied to the functions dependent on $r_{2}$ and $r_{3}$. For example,
@@
$\mathcal{F}_{\lambda}(r,R)$ $=\frac{\hat{\lambda}^{2}}{2}\int_{-1}^{1}dwP_{\lambda}(w)Y\!\left(\sqrt{M_{2}^{2}r^{2}+R^{2}-2M_{2}rRw}\right),$ (108)

### v2

@@ II.3 Decuplet saturation approximation
Here $\Delta$ denotes the average decuplet–octet baryon mass splitting, and we take $\Delta=203$ MeV, which is evaluated using the masses of the baryons considered, namely, the proton, neutron, $\Delta(1232)$, $\Xi^{0}$, $\Xi^{-}$, and $\Xi(1530)$ [32]. We employ the large-$N_{c}$ value for the LEC $C$, namely, $C=3g_{A}/4$} [35, 23, 44]. Then, under the DSA, the LECs $G_{e}^{(\sigma)}$ and $G_{f}^{(\sigma)}$, which enter $U_{\mathrm{OPE}}^{(\sigma)}$, become independent of $\sigma$:
@@ II.4 Correlation function
where $F_{L}$ is the regular Coulomb wave function with the orbital angular momentum $L$ and $q$ is the relative momentum between $d$ and $\Xi^{-}$. The source function $\mathcal{S}$ used here is the same as that employed in Ref. [33]:
@@
$V_{\mathrm{C}}(R)=\begin{dcases}-\frac{e^{2}}{2R_{\mathrm{C}}}\left(3-\frac{R^{2}}{R_{\mathrm{C}}^{2}}\right)&(R\leq R_{\mathrm{C}})\\ -\frac{e^{2}}{R_{\mathrm{C}}}&(R>R_{\mathrm{C}}),\end{dcases}$ (77)
@@ III.1 Comparison of potentials
Figure 6(a) shows $U_{\mathrm{TPE}}^{(u_{3})}$, $U_{\mathrm{OPE}}^{(\sigma)}$, and $U_{\mathrm{ct}}^{(\sigma)}$ as a function of $R$, and is depicted by the solid, dashed, and dotted lines, respectively. Note that $U_{\mathrm{TPE}}^{(u_{1})}$ defined by Eq. (47) becomes zero under the DSA. As a representative case, we fix both $G^{\prime}=1$} and $H^{\prime}=1$}, with the cutoff $R_{0}=1.0$} fm. One sees that the contact term can have the largest magnitude at $R\lesssim 1.0$} fm. Beyond this range, the TPE and OPE terms become dominant although their absolute magnitude is much smaller compared with that of the contact term near the origin.
In Fig. (6)(b), we compare $U_{\mathrm{3BF}}^{(\sigma)}$ with $U_{\mathrm{2BF}}^{(\sigma)}$. In this case, we select two sets of the LECs, one giving the most attractive and the other the most repulsive potential, in order to maximize the variation in $U_{\mathrm{3BF}}^{(\sigma)}$. These parameter sets correspond to the solid and dashed lines, with $(G^{\prime},H^{\prime})=(1.0,-1.0)$} and $(G^{\prime},H^{\prime})=(-1.0,1.0)$}, respectively. The dotted and dash-dotted lines are $U_{\mathrm{2BF}}^{(\sigma)}$ with $\sigma=1/2$ and $3/2$, respectively. In both sets of the LECs, $U_{\mathrm{3BF}}^{(\sigma)}$ are found to be weaker in magnitude and shorter in range compared with $U_{\mathrm{2BF}}^{(\sigma)}$.
@@ III.2 Effect of $\Xi NN$ three-baryon force on correlation function
Figure 7(a) represents the calculated result of the correlation function $C_{d\Xi^{-}}$ as a function of $q$. We vary the LECs within the range $1\leq G^{\prime}\leq 1$ and $1\leq H^{\prime}\leq 1$, and thus, the most attractive and most repulsive cases are taken as the boundaries, forming a band. The solid line corresponds to the correlation function calculated without the $\Xi NN$ 3BF. The cutoff and source size are set to be $R_{0}=1.0$} fm and b=1.2{b=1.2} fm, respectively. To quantify the effect of the 3BF, we present in Fig. 7(b) the maximum relative deviation of the band boundaries from the central (solid) line at each $q$. As seen in these two figures, the effect of the 3BF on $C_{d\Xi^{-}}$ are very limited, at most around $4\%$. This is even smaller than the effect of deuteron breakup, which was found to be about $6\%$$8\%$ [33]. One possible reason for the small 3BF effect on the correlation function is the short-range nature of the 3BFs. As shown in Fig. 6, $U_{\mathrm{3BF}}^{(\sigma)}$ becomes nearly zero around $R\gtrsim 1.0$ fm, in contrast to $U_{\mathrm{2BF}}^{(\sigma)}$. We numerically confirmed that $C_{d\Xi^{-}}$ is mainly sensitive to the potential in the range $1.0\lesssim R\lesssim 2.0$ fm, although it depends on $b$.
@@ III.3 Variation of parameters
To conclude this section, we consider whether the loosely bound nature of the deuteron can explain the small effect of the $\Xi NN$ 3BF on $C_{d\Xi^{-}}$. To examine this, we artificially change the deuteron binding energy $E_{\mathrm{BE}}$ by modulating the strength of the Argonne $V4^{\prime}$ potential. Figures 10(a)–(c) show $U_{\mathrm{2BF}}^{(\sigma)}$ and $U_{\mathrm{3BF}}^{(\sigma)}$ computed with $E_{\mathrm{BE}}/\mathrm{nucleon}=1.12$$2.54$, and $6.98$ MeV, corresponding to that of the realistic deuteron, triton and $\alpha$ particle, respectively. Even though $E_{\mathrm{BE}}$ is increased, the contribution of $U_{\mathrm{3BF}}^{(\sigma)}$ relative to $U_{\mathrm{2BF}}^{(\sigma)}$ does not change significantly. Figures 10(d)–(f) display $C_{d\Xi^{-}}$ computed with the potentials in Figs. 10(a)–(c), respectively. With increasing $E_{\mathrm{BE}}$, the magnitudes of the potentials grow, enhancing $C_{d\Xi^{-}}$. This enhancement mainly originates from $U_{\mathrm{2BF}}^{(\sigma)}$ rather than $U_{\mathrm{3BF}}^{(\sigma)}$. Indeed, even for the strongly bound case in Fig. 10(f), the effect of the $\Xi NN$ 3BF remains below $8\%$. Therefore, we conclude that the weak binding of the deuteron is not the primary reason for the small effect of the 3BF on the correlation function.
IV Summary
@@ Acknowledgements.
The authors thank T. Hyodo, A. Jinno, S. Ogawa, and Y. Chazono for valuable comments and discussions. This work was supported by JST ERATO Grant No. JPMJER2304, as well as JSPS KAKENHI Grants No. JP21K13919 and No. JP23KK0250.
@@ Appendix A SU(3) COEFFICIENTS OF VERTICES RELEVANT TO $\Xi NN$ THREE-BARYON FORCE
Table 1: SU(3) coefficients $N_{BB\phi}$ relevant to the $\Xi NN$-TPE and OPE potentials. The axial vector coupling coefficients $g_{A}$ and $g_{B}$ are defined in Sec. II.1.2.
@@
Table 4: Lists of the contact coefficients $N_{\begin{subarray}{c}BBB\\ BBB\end{subarray}}^{a}$ defined by $N_{\begin{subarray}{c}BBB\\ BBB\end{subarray}}^{a}=\sum_{f=1}^{11}\tilde{t}^{f,a}N_{\begin{subarray}{c}BBB\\ BBB\end{subarray}}^{f,a}$, which enter the $\Xi NN$-contact potentials as in Eqs. (33). The coefficients are all written in terms of the LECs $C_{i}$ introduced in Sec. II.1.4.
@@ B.1 Central components of $\Xi NN$ potential
$X_{123(0)}^{456}$ $=\dfrac{g_{B}d}{6f_{\pi}^{2}}\dfrac{q_{1}^{2}}{q_{1}^{2}+m_{\pi}^{2}}\left[(\bm{\tau}_{2}-\bm{\tau}_{3})\cdot\bm{\tau}_{1}\,(\bm{\sigma}_{2}-\bm{\sigma}_{3})\cdot\bm{\sigma}_{1}+(\bm{\tau}_{1}\times\bm{\tau}_{2})\cdot\bm{\tau}_{3}\,(\bm{\sigma}_{1}\times\bm{\sigma}_{2})\cdot\bm{\sigma}_{3}\right],$ (84)
$X_{231(0)}^{564}$ $=\dfrac{g_{A}}{6f_{\pi}^{2}}\dfrac{q_{2}^{2}}{q_{2}^{2}+m_{\pi}^{2}}\left[(e_{1}\bm{\sigma}_{3}+e_{2}\bm{\sigma}_{1})\cdot\bm{\sigma}_{2}\,\bm{\tau}_{2}\cdot\bm{\tau}_{3}+(e_{3}\bm{\sigma}_{3}+e_{4}\bm{\sigma}_{1})\cdot\bm{\sigma}_{2}\,\bm{\tau}_{1}\cdot\bm{\tau}_{2}-e_{5}(\bm{\tau}_{1}\times\bm{\tau}_{2})\cdot\bm{\tau}_{3}\,(\bm{\sigma}_{1}\times\bm{\sigma}_{2})\cdot\bm{\sigma}_{3}\right],$ (85)
$X_{312(0)}^{645}$ $=\dfrac{g_{A}}{6f_{\pi}^{2}}\dfrac{q_{3}^{2}}{q_{3}^{2}+m_{\pi}^{2}}\left[(e_{4}\bm{\sigma}_{1}+e_{3}\bm{\sigma}_{2})\cdot\bm{\sigma}_{3}\,\bm{\tau}_{3}\cdot\bm{\tau}_{1}+(e_{2}\bm{\sigma}_{1}+e_{1}\bm{\sigma}_{2})\cdot\bm{\sigma}_{3}\,\bm{\tau}_{2}\cdot\bm{\tau}_{3}-e_{5}(\bm{\tau}_{1}\times\bm{\tau}_{2})\cdot\bm{\tau}_{3}\,(\bm{\sigma}_{1}\times\bm{\sigma}_{2})\cdot\bm{\sigma}_{3}\right],$ (86)
$X_{312(0)}^{564}$ $=\dfrac{g_{A}}{6f_{\pi}^{2}}\dfrac{q_{3}^{2}}{q_{3}^{2}+m_{\pi}^{2}}\left[\{f_{1}\bm{\sigma}_{1}+f_{2}\bm{\sigma}_{2}+f_{3}i(\bm{\sigma}_{1}\times\bm{\sigma}_{2})\}\cdot\bm{\sigma}_{3}\,\bm{\tau}_{3}\cdot\bm{\tau}_{1}+\{f_{2}\bm{\sigma}_{1}+f_{1}\bm{\sigma}_{2}+f_{3}i(\bm{\sigma}_{1}\times\bm{\sigma}_{2})\}\cdot\bm{\sigma}_{3}\,\bm{\tau}_{2}\cdot\bm{\tau}_{3}\right$.
@@
$X_{231(0)}^{645}$ $=\dfrac{g_{A}}{6f_{\pi}^{2}}\dfrac{q_{2}^{2}}{q_{2}^{2}+m_{\pi}^{2}}\left[\{f_{1}\bm{\sigma}_{3}+f_{2}\bm{\sigma}_{1}-f_{3}i(\bm{\sigma}_{3}\times\bm{\sigma}_{1})\}\cdot\bm{\sigma}_{2}\,\bm{\tau}_{2}\cdot\bm{\tau}_{3}+\{f_{2}\bm{\sigma}_{3}+f_{1}\bm{\sigma}_{1}-f_{3}i(\bm{\sigma}_{3}\times\bm{\sigma}_{1})\}\cdot\bm{\sigma}_{2}\,\bm{\tau}_{1}\cdot\bm{\tau}_{2}\right$.
@@ B.2 Matrix elements of spin–isospin operators
with $\mu=\pm 1$. In Eq. (90), $\sigma$ and $m_{\sigma}$ represent the total spin of the system and its third component, respectively. The isospin and its third component of the system are fixed as $1/2$ and 1/2-1/2, respectively. Using these states, we evaluate the matrix elements of all spin and isospin operators involved in the potentials given in Sec. B.1.
Hereafter we redefine the particle indices $\alpha$ as $\Xi\to 1$, $p\to 2$, and $n\to 3$. The matrix elements of the spin operators, $\bm{\sigma}_{1}\cdot\bm{\sigma}_{2}$ and $\bm{\sigma}_{1}\cdot\bm{\sigma}_{3}$, depend on $\sigma$:
$\Braket{\Upsilon_{\sigma m_{\sigma}}|\bm{\sigma}_{1}\cdot\bm{\sigma}_{2}|\Upsilon_{\sigma m_{\sigma}}}$ $=\Braket{\Upsilon_{\sigma m_{\sigma}}|\bm{\sigma}_{3}\cdot\bm{\sigma}_{1}|\Upsilon_{\sigma m_{\sigma}}}=\begin{dcases}-2&(\sigma=1/2)\\ 1&(\sigma=3/2).\end{dcases}$ (93)
The other spin operators yield $\sigma$-independent expressions:
@@ B.3 Fourier transform and multipole expansion
where the functions $Y$ and $Z$ are defined by Eqs. (54) and (56), respectively. The arguments of the spherical harmonics $Y_{\lambda\mu}$ ($\lambda\leq\mu\leq\lambda$) are defined by $\hat{\bm{r}}_{i}=\bm{r}_{i}/r_{i}$. To carry out the above calculations, the following relation is used [29]:
@@
$W^{(\sigma)}(\bm{r},\bm{R})$ $=W_{\mathrm{TPE}}(\bm{r},\bm{R})+W_{\mathrm{OPE}}^{(\sigma)}(\bm{r},\bm{R})+W_{\mathrm{ct}}^{(\sigma)}(\bm{r},\bm{R}),$ (103)
$W_{\mathrm{TPE}}(\bm{r},\bm{R})$ $=-\frac{g_{A}^{2}m_{\pi}^{6}u_{1}}{4\sqrt{3}\pi f_{\pi}^{4}}Z_{1}(m_{\pi}r_{2})Z_{1}(m_{\pi}r_{3})\left[Y_{1}\left(\hat{\bm{r}}_{2}\right)\otimes Y_{1}\left(\hat{\bm{r}}_{3}\right)\right]_{00}$
$-\dfrac{g_{A}^{2}m_{\pi}^{6}u_{3}}{12\pi f_{\pi}^{4}}\left[\dfrac{1}{8\pi}Y(m_{\pi}r_{2})Y(m_{\pi}r_{3})+\dfrac{1}{\sqrt{5}}Z(m_{\pi}r_{2})Z(m_{\pi}r_{3})\left[Y_{2}\left(\hat{\bm{r}}_{2}\right)\otimes Y_{3}\left(\hat{\bm{r}}_{2}\right)\right]_{00}\right],$ (104)
@@
where $\bm{r}_{2}$ and $\bm{r}_{3}$ can be expressed as functions of $\bm{r}$ and $\bm{R}$ by Eqs. (43) and (44). The function $Z_{1}$, as well as the constants $G_{e}^{(\sigma)}$, $G_{f}^{(\sigma)}$, and $H^{(\sigma)}$, are all defined in Sec. II.2.
Then, by applying the regularization described in Sec. II.2 to $W^{(\sigma)}$, we can perform the integration over $\bm{r}$ and $\hat{\bm{R}}$ in Eq. (45), for which the multipole expansion is applied to the functions dependent on $r_{2}$ and $r_{3}$. For example,
@@
$\mathcal{F}_{\lambda}(r,R)$ $=\frac{\hat{\lambda}^{2}}{2}\int_{-1}^{1}dwP_{\lambda}(w)Y\!\left(m_{\pi}\sqrt{M_{2}^{2}r^{2}+R^{2}-2M_{2}rRw}\right),$ (108)